OpenStudy (anonymous):

Last week, the ratio of laptops to desktops sold at a computer store was 2:3. Eighteen desktop models were sold. How many laptop models were sold?

OpenStudy (anonymous):

2/3 = x/13

OpenStudy (anonymous):

*18

OpenStudy (anonymous):

l : d = 2:3 \[\frac{l}{d}=\frac{2}{3}\] let the common parameter be x. \[l=2x\] \[d=3x\] it is given that l+d=18 so, 2x+3x=18 find x. then find l(=2x)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

it is given d=3x=18 therefore , x=6 therefore, l=2x=12